OpenStudy (anonymous):

write the fraction in lowest terms 45c^2/81cd

OpenStudy (lalaly):

\[\large{\frac{45c^2}{81cd}}=\frac{5c}{9d}\]

OpenStudy (anonymous):

5c/9d